# On modules whose dual is of finite Gorenstein dimension: The $k$-torsion, homological criteria, and applications in modules of differentials of order $n$

### v4

@@ Abstract.
In this paper, we aim to obtain some results under the condition that the dual of a module has finite Gorenstein dimension. In this direction we derive results involving vanishing of Ext as well as the freeness or totally reflexivity of modules. For instance, we provide a generalization of a celebrated theorem by Evans and Griffith, obtain criteria for the totally reflexivity of modules over Cohen-Macaulay rings as well as of locally totally reflexive modules on the punctured spectrum, and recover a result by Araya. Moreover, we prove that the Auslander-Reiten conjecture holds true for all finitely generated modules $M$ over a Noetherian local ring $R$ such that $\operatorname{G-dim}_{R}(\operatorname{Hom}_{R}(M,R))<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(M,M))<\infty$. Additionally, we derive Gorenstein criteria under the condition that the dual of certain modules is of finite Gorenstein dimension. Furthermore, we explore some applications in the theory of the modules of Kähler differentials of order $n\geq 1$, specifically concerning the $k$-torsionfreeness of these modules and the Herzog-Vasconcelos’s conjecture.
@@ 1. Introduction
The Gorenstein dimension is a well-developed topic in commutative algebra, as evidenced by various works such as [5, 14, 44, 50, 18, 52]. Despite the extensive research in this area, there are relatively few results concerning the condition of the Gorenstein dimension of the dual of a certains modules being finite. The aim of this work is to obtain results in this direction, including some that involve vanishing of Ext and homological criteria.
We prove the following theorem, which is a generalization of a result of Evans and Griffith ([15, Theorem 3.8]) and is an analogous version of [14, Proposion 2.4] when $C=R$.
1.1 Theorem.
Let $R$ be a Noetherian local ring and $k$ be a non-negative integer. Let $M$ be a finitely generated non-zero $R$-module such that $M^{\ast}$ has locally finite Gorenstein dimension on $\widetilde{X}^{k-1}(R)$. Then:
@@
This theorem provides an answer to [37, Question 1.1]. Additionally, it allows us to recover one of the results described in the main theorem of [52].
1.2 Corollary ([52, Theorem 5.8(1)]).
Let $R$ be a Noetherian local ring, $M$ be a finitely generated $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and $k$ be a non-negative integer. If $M$ satisfies $(\widetilde{S}_{k})$, then $M$ is $k$-torsionfree.
As an application of Corollary 1.2 and [50, Theorem 42], we characterize under certain conditions and in terms of vanishing of Ext modules, the total reflexivity of a finitely generated $R$-module $M$ with $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. We obtain the following.
1.3 Theorem.
Let $R$ be a Noetherian local ring of depth $t$, and let $M$ be a finitely generated $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then:
@@
    Let $R$ be a Cohen-Macaulay local ring and $M$ is non-zero. If $n$ is a non-negative integer such that $\operatorname{depth}_{R}(M)\geq n$ and $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq t-n,$ then $M$ is totally reflexive.
@@
Additionally, part of this work is motivated by the celebrated Auslander-Reiten conjecture ([6]), which states the following.
1.4 Conjecture.
(Auslander-Reiten) Let $R$ be a Noetherian local ring and let $M$ be a finitely generated $R$-module. If $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{i}(M,M)=0$ for all $i>0$, then $M$ is free.
@@
This conjecture is still open but is true in several cases. For a list of some of them we refer the reader to [30], it introduction and to the recent work [12]. In [2], Araya proved the following result that provided an positive answer to Auslander-Reiten conjecture, and that implies that it is valid over normal Gorenstein local rings of dimension at least two.
1.5 Theorem ([2, Corollary 10]).
@@
Let $R$ be a Gorenstein local ring of dimension $d\geq 2$. Let $M$ be a maximal Cohen-Macaulay $R$-module. If $M$ is locally free on the punctured spectrum of $R$ and $\operatorname{Ext}_{R}^{d-1}(M,M)=0$, then $M$ is free.
@@
Subsequently, results concerning the context of Gorenstein rings that recover this theorem have appeared. For instance, the last theorem in [42, Section 1] and [3, Corollary 1.6]. It is important to mention that in a more general context, [30, Proposition 2.10(4)] and [8, Proposition 1.7] are improvements of these results. In terms of the finiteness of the Gorenstein dimension of the dual of a module, we achieve the following results that recover the last theorem in [42, Section 1], [3, Corollary 1.6] and of course Theorem 1.5.
1.6 Theorem.
@@
Let $R$ be a Noetherian local ring of depth $t\geq 1$ and let $M$ be a finitely generated $R$-module. Suppose that $M$ is locally free on the spectrum punctured such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. If $0pt(M)\geq t$ and $\operatorname{Ext}_{R}^{t-1}(M,M)=0$, then $M$ is free.
1.7 Theorem.
Let $R$ be a Noetherian ring satisfying $(S_{1})$ and $X$ a subset of $\operatorname{Spec}(R)$ containing $X^{1}(R)$. Let
$s:=\inf\{\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\mid\mathfrak{p}\in\operatorname{Spec}(R)\backslash X\}$ and $t:=\sup\{\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\mid\mathfrak{p}\in\operatorname{Spec}(R)\backslash X\}$.
@@
Let $M$ be a finitely generated $R$-module and suppose that $M$ is locally free on $X$ and that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. If $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq i\leq t$ and $s-1\leq j\leq t-1$, then $M$ is projective.
@@
On the other hand, in [12, Corollary 6.9(2)], Dey and Ghosh proved that the Auslander-Reiten conjecture holds for finitely generated $R$-modules $M$ such that $\operatorname{G-dim}_{R}(M)<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(M,M))<\infty$. Motivated by this, we analyze if the conjecture holds if we consider $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ instead of $\operatorname{G-dim}_{R}(M)<\infty$. Consequently, we get the following result that affirms that it holds.
1.8 Theorem.
@@
Let $R$ be a Noetherian local ring of depth $t$, and let $M$ be a finitely generated $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then the following conditions are equivalent:
@@
    $\operatorname{Hom}_{R}(M,M)$ is free and $\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $11\leq j\leq t-$ 1.
@@
    $\operatorname{Hom}_{R}(M,M)$ has finite projective dimension and $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq i\leq t$ and $11\leq j\leq t-$ 1.
Additionally, we also derive some Gorenstein criteria involving the condition that the dual of certain $R$-modules are of finite Gorenstein dimension. We do this essentially by discussing of some questions about that. We obtain the following.
1.9 Theorem.
Let $R$ be a Noetherian local ring.Then $R$ is Gorenstein in each one of the following cases:
@@
    $R$ be a Cohen-Macaulay local ring and there exists an $R$-module $M$ such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and it satisfies one of the following:
@@
    $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ for all finitely generated $R$-modules $M$ such that $M^{\ast}\not=0$.
@@
    $R$ has positive depth, and every finitely generated $R$-modules $M$ whose dual $M^{\ast}$ is non-zero and of finite Gorenstein dimension, is also of finite Gorenstein dimension.
@@
In the last part of this paper, we focus on the Kähler differential modules $\Omega_{X/Y}^{(n)}$ and the derivation modules $\mathcal{D}er_{Y}^{n}(X)$ defined over locally Noetherian schemes and analytic spaces, where $n\geq 1$ is an integer. The motivation for studying these modules arises from the Zariski-Lipman conjecture [33], which states: Let $X$ be a complex variety such that the tangent sheaf $\mathcal{T}_{X}:=\mathcal{H}om_{\mathcal{O}_{X}}(\Omega_{X}^{(1)},\mathcal{O}_{X})$ is locally free. Then $X$ is smooth. This conjecture, proposed about 50 years ago, is still open, although it has been resolved in some special cases, see, for instance, [39, Section 4]. Following the same direction as the Zariski-Lipman Conjecture, there is also a homological version of it, independently proposed by Herzog and Vasconcelos, which predicts that if ${\rm pd}_{R}{\rm Der}_{k}(R)\,<\,\infty$, then ${\rm Der}_{k}(R)$ is free. Unlike the former, this problem seems to be widely open (with exceptions in specific cases; see, for example, [26, Section 4] and [40]). There is also another conjecture called the Strong Zariski-Lipman Conjecture: If ${\rm pd}_{R}{\rm Der}_{k}(R)<\infty$, then $R$ is regular. Due to these conjectures, the following questions arise:
1.10 Questions.
  • (i)
    (Generalizations of Herzog-Vasconcelos’s conjecture (GHVC)) For some integer $n\geq 1$, we have ${\rm pd}_{\mathcal{O}_{X}}{\mathcal{D}er}_{Y}^{n}(X)\,<\,\infty$. Under what assumptions on $X$ and $Y$, and for which values of $n$, does this imply that ${\mathcal{D}er}_{Y}^{n}(X)$ is locally free?
  • (ii)
    (Strong generalizations of Herzog-Vasconcelos’s conjecture (SGHVC)) For some integer $n\geq 1$, we have ${\rm pd}_{\mathcal{O}_{X}}{\mathcal{D}er}_{Y}^{n}(X)\,<\,\infty$. Under what assumptions on $X$ and $Y$, and for which values of $n$, does this imply that $X$ is locally smooth?
It is worth noting that these questions and others appear in the works of Graf [22] (in the case of algebraic varieties and analytic spaces) and in the case of affine schemes in [35]. Historically, to answer these questions, it is necessary to study the concepts of torsion, cotorsion, torsion-free, and reflexive of the Kähler differential module (see, for example, a summary in [39, p. 133] and [22]). For this reason,in the Sections 8 and 9, we obtain some direct applications connecting these concepts with the concepts of $k$-torsion-free and $k$-syzygy. Consequently, we also obtain some special answers for GHVC and SGHVC.
@@
The organization of this paper is as follows. In Section 2 we provide definitions, notations, and some results that are considered in this paper. In Section 3 we prove Theorem 1.1 and explore some of its consequences. In Section 4, we prove Theorem 1.3, and in Section 5 we provide freeness criteria including Theorem 1.7 and 1.8. In Section 6, we discuss some questions about of the Gorensteiness of a ring involving Gorenstein dimension of the dual of modules, and prove Theorem 1.9. In Section 7, we introduce notation and review established results regarding Kähler differential modules and differential modules. Then, in Section 8, we investigate the properties of the Kähler differential module $\Omega_{X/Y}^{(n)}$, examining its local freeness, reflexivity, $k$-torsion-freeness, and $k$-syzygy properties, specially when $k=1,2$. Finally, in the last section, we provide a partial response to the SGHV and SGHV conjectures.
@@ 2. Setup and preliminaries
Throughout this paper, $R$ is a Noetherian ring, and all $R$-modules are considered to be finitely generated. Whenever $R$ is local, we denote by $\mathfrak{m}$ and $k$ its maximal ideal and its residual field respectively.
@@
Let $M,N$ be two $R$-modules, and $k$ be a non-negative integer.
@@
    Let
    $\boldsymbol{P}:\cdots\longrightarrow P_{i}\stackrel{{\scriptstyle\varphi_{i}}}{{\longrightarrow}}P_{i-1}\longrightarrow\cdots\longrightarrow P_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}P_{0}\stackrel{{\scriptstyle\varphi_{0}}}{{\longrightarrow}}M\longrightarrow 0$
    be a projective resolution of resolution of $M$.
@@
      The Auslander Transpose of $M$, denoted by $\operatorname{Tr}(M)$, is defined as the cokernel of the induced map $\varphi_{1}^{\ast}:P_{0}^{\ast}\to P_{1}^{\ast}.$ When $k\geq 1$, we set
      $\mathcal{T}_{k}(M)=\operatorname{Tr}(\Omega^{k-1}(M)).$
@@
    $M$ is a $k$-syzygy if there exists an exact sequence
    $0\rightarrow M\rightarrow P_{k-1}\rightarrow P_{k-2}\rightarrow\cdots\rightarrow P_{1}\rightarrow P_{0},$
    where each $P_{i}$ is of projective dimension.
@@
    We say that $M$ satisfies $(S_{k})$ (resp. $(\widetilde{S}_{k})$) if $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\min\{k,\operatorname{ht}(\mathfrak{p})\}$ (resp. $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\min\{k,0pt_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\})$.
@@
The $R$-modules $\operatorname{Tr}(M)$ and $\Omega^{k}(M)$ ($k\geq 0$) are uniquely determined up to projective summands. It is easy to see that $\operatorname{Tr}(\operatorname{Tr}(M))\approx M$. Note that $M$ is a $k$-syzygy if and only if there exists an $R$-module $L$ such that $M\approx\Omega^{k}(L)$. Besides, for every $R$-module $N$, there exists an exact sequence of functors (see [5, Theorem 2.8]).
(2.1) $0\rightarrow\operatorname{Ext}_{R}^{1}\left(\mathcal{T}_{k+1}(M),-\right)\rightarrow\operatorname{Tor}_{k}^{R}(M,-)\rightarrow\operatorname{Hom}_{R}(\operatorname{Ext}_{R}^{k}(M,R),-)\rightarrow\operatorname{Ext}_{R}^{2}\left(\mathcal{T}_{k+1}(M),-\right).$
Whenever $R$ is local, we consider the Auslander tranpose and the syzygies (of an $R$-module) defined using minimal free resolutions. In this case, these $R$-modules are defined uniquely up to isomorphism (rather than projective equivalence).
The notion of Gorenstein dimension was introduced by Auslander [36] and developed by Auslander and Bridger in [5].
@@
    The Gorenstein dimension of $M$, denoted by $\operatorname{G-dim}_{R}(M)$, is defined to be the infimum of all non-negative integers $k$ such that there exists an exact sequence
    $0\rightarrow G_{k}\rightarrow\cdots\rightarrow G_{0}\rightarrow M\rightarrow 0,$
    where each $G_{i}$ is totally reflexive.
@@
    ([9, Theorem 1.2.7]) If $M\not=0$ and $\operatorname{G-dim}_{R}(M)<\infty$, then
    $\operatorname{G-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,R)\not=0\}.$
@@
    ([5, Lemma 3.19(1)]) $M$ is totally reflexive if and only if $\operatorname{Tr}_{R}(M)$ is totally reflexive.
@@
    ([9, Corollary 1.2.9]) Let $0\to M\to N\to L\to 0$ be an exact sequence of $R$-modules. If two $R$-modules of the sequence have finite Gorenstein dimension, then so has the third.
@@
    ([9, Proposition 1.2.10]) $\operatorname{G-dim}_{R}(M)\leq\operatorname{pd}_{R}(M)$ with equality if $\operatorname{pd}_{R}(M)<\infty$.
@@
    ([9, Theorem 1.4.8]) If $M\not=0$, then $\operatorname{G-dim}_{R}(M)<\infty$, then
    $\operatorname{G-dim}_{R}(M)+\operatorname{depth}_{R}(M)=\operatorname{depth}_{R}(R).$
    In particular, $\operatorname{G-dim}_{R}(M)<\infty$.
@@
    ([9, Theorem 1.4.9]) The following conditions are equivalent:
@@
      $\operatorname{G-dim}_{R}(M)<\infty$ for all (finitely) $R$-modules $M$.
@@
The formula given in Fact 2.3(5) is known as the Auslander-Bridger formula. From the definition of the Auslander transpose and Fact 2.3(3) we can derive the following remark.
@@
Let $R$ be a ring and let $M$ be an $R$-module. Then $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ if and only if $\operatorname{G-dim}_{R}(\operatorname{Tr}_{R}(M))<\infty$.
3. A generalization of a theorem of Evans and Griffith
The main goal of this section is to provide a generalization of a celebrated result by Evans and Griffith ([15, Theorem 3.8]), which characterizes a module to be $k$-torsionfree, $k$-syzygy, involving the condition $(\widetilde{S}_{k})$.

### v5

@@ Abstract.
In this paper, we aim to obtain some results under the condition that the dual of a module over a commutative Noetherian ring has finite Gorenstein dimension. In this direction, we derive results involving vanishing of Ext as well as the freeness or totally reflexivity of modules. For instance, we provide a generalization of a celebrated theorem by Auslander and Bridger, obtain criteria for the totally reflexivity of modules over Cohen-Macaulay rings as well as of locally totally reflexive modules on the punctured spectrum, and recover a result by Araya. Moreover, we prove that the Auslander-Reiten conjecture holds true for all finitely generated modules $M$ over a commutative Noetherian ring $R$ such that $\operatorname{G-dim}_{R}(\operatorname{Hom}_{R}(M,R))<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(M,M))<\infty$. Additionally, we derive Gorenstein criteria under the condition that the dual of certain modules is of finite Gorenstein dimension. Furthermore, we explore some applications in the theory of the modules of Kähler differentials of order $n\geq 1$, specifically concerning the $k$-torsionfreeness of these modules and the Herzog-Vasconcelos conjecture.
Key words and phrases:
Gorenstein dimension, Algebraic dual, $k$-torsion, Totally reflexive modules, Free modules, Vanishing of Ext, Auslander-Reiten conjecture, Gorenstein ring, Module of Kähler differentials, Herzog-Vasconcelos conjecture
@@ 1. Introduction
The Gorenstein dimension is a well-developed topic in commutative algebra, as evidenced by various works such as [Stablemoduletheory, LinkageOfModulesAndTheSerreConditions, ktorsionlessmoduleswithfiniteGorensteindimension, GorensteinDimensionAndTorsionOfModulesOverCommutativeNoetherianRings, GorensteinringsviahomologicaldimensionsandsymmetryinvanishingofExtandTatecohomology, HomologicalDimensionsOfRigidModules]. Despite the extensive research in this area, there are relatively few results concerning the condition of the Gorenstein dimension of the dual of certain modules being finite. The aim of this work is to obtain results in this direction, including some that involve vanishing of Ext and homological criteria.
Let $R$ be a (commutative) Noetherian ring. We prove the following theorem, which is a generalization of a result that Evans and Griffith attribute to Auslander and Bridger in [Syzygies, Theorem 3.8] (see also [AscentOfFiniteCohen–MacaulayType, Lemma 1.3]) and is an analogous version of [LinkageOfModulesAndTheSerreConditions, Proposition 2.4] when $C=R$.
1.1 Theorem (See Theorem 3.2).
Let $k$ be a non-negative integer. Let $M$ be a finitely generated $R$-module such that $M^{\ast}$ has locally finite Gorenstein dimension on $\widetilde{X}^{k-1}(R)$. Then:
@@
This theorem provides an answer to [WhenArenSysymodulesnTorsionfree, Question 1.1]. Additionally, it allows us to recover one of the results described in the main theorem of [HomologicalDimensionsOfRigidModules].
1.2 Corollary ([HomologicalDimensionsOfRigidModules, Theorem 5.8(1)]).
Suppose that $R$ is local. Let $M$ be a finitely generated $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and $k$ be a non-negative integer. If $M$ satisfies $(\widetilde{S}_{k})$, then $M$ is $k$-torsionfree.
As an application of Corollary 1.2 and [GorensteinDimensionAndTorsionOfModulesOverCommutativeNoetherianRings, Theorem 42], we characterize the totally reflexivity of a module whose dual is of finite Gorenstein dimension in terms of certain Ext vanishing and some other conditions. We obtain the following.
1.3 Theorem (See Theorems 4.1, 4.8 and 4.9).
Suppose that $R$ is local of depth $t$. Let $M$ be a finitely generated $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then:
@@
    Let $R$ be a Cohen-Macaulay local ring. If $n$ is a non-negative integer such that $\operatorname{depth}_{R}(M)\geq n$ and $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq t-n,$ then $M$ is totally reflexive.
@@
Additionally, part of this work is motivated by the celebrated Auslander-Reiten conjecture ([OnageneralizedversionoftheNakayamaconjecture]), which states the following.
1.4 Conjecture (Auslander-Reiten).
Let $M$ be a finitely generated $R$-module. If $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{i}(M,M)=0$ for all $i>0$, then $M$ is projective.
@@
This conjecture is still open but it is true in several cases. For a list of some of them, we refer the reader to the introductions of [injectivedimensiontakahashi, AuslanderReitenConjectureforNormalRings, FiniteHomologicalDimensionOfHomv3]. In [TheAuslander-ReitenconjectureforGorensteinrings], Araya proved the following result that provided a positive answer to Auslander-Reiten conjecture, and that implies that it is valid for normal Gorenstein local rings of dimension at least two.
1.5 Theorem ([TheAuslander-ReitenconjectureforGorensteinrings, Corollary 10]).
@@
Suppose that $R$ is Gorenstein local of dimension $d\geq 2$. Let $M$ be a maximal Cohen-Macaulay $R$-module. If $M$ is locally free on the punctured spectrum of $R$ and $\operatorname{Ext}_{R}^{d-1}(M,M)=0$, then $M$ is free.
@@
Subsequently, results concerning the context of Gorenstein rings that recover this theorem have appeared. For instance, the last theorem in [AnAuslander–Reitenprincipleinderivedcategories, Section 1] and [OnthevanishingofselfextensionsoverCohen-Macaulaylocalrings, Corollary 1.6]. It is important to mention that in a more general context, [AuslanderReitenConjectureforNormalRings, Proposition 2.10(4)] and [TwoTheoremsOnTheVanishingOfExt, Proposition 1.7] are improvements of these results. In terms of the finiteness of the Gorenstein dimension of the dual of a module, we achieve the following results that recover the last theorem in [AnAuslander–Reitenprincipleinderivedcategories, Section 1], [OnthevanishingofselfextensionsoverCohen-Macaulaylocalrings, Corollary 1.6] and, of course, Theorem 1.5.
1.6 Theorem (See Corollary 5.6).
@@
Suppose that $R$ is local. Let $M$ be a finitely generated $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Suppose that $M$ is locally free on the punctured spectrum of $R$. If $0pt(M)\geq t$ and $\operatorname{Ext}_{R}^{t-1}(M,M)=0$, then $M$ is free.
1.7 Theorem (See Theorem 5.7).
Let $R$ satisfy $(S_{1})$ and let $X$ be a subset of $\operatorname{Spec}(R)$ containing $X^{1}(R)$. Let
[…]
@@
Let $M$ be a finitely generated $R$-module. Suppose that $M$ is locally free on $X$ and that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. If $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq i\leq t$ and $s-1\leq j\leq t-1$, then $M$ is projective.
@@
On the other hand, in [FiniteHomologicalDimensionOfHomv3, Corollary 6.9(2)], Dey and Ghosh proved that the Auslander-Reiten conjecture holds for finitely generated modules $M$ over a (commutative) Noetherian ring $R$ such that $\operatorname{G-dim}_{R}(M)<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(M,M))<\infty$. Motivated by this, we analyze if this fact holds if we consider $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ instead of $\operatorname{G-dim}_{R}(M)<\infty$. Consequently, we get the following result, which not only affirms that this holds but also strengthens [injectivedimensiontakahashi, Corollary 2.14].
1.8 Theorem (See Theorem 5.11).
@@
Suppose that $R$ is local of depth $t$. Let $M$ be a finitely generated $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then the following conditions are equivalent:
@@
    $\operatorname{Hom}_{R}(M,M)$ is free and $\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq j\leq t-1$.
@@
    $\operatorname{Hom}_{R}(M,M)$ has finite projective dimension, and $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq i\leq t$ and $1\leq j\leq t-1$.
Additionally, we also derive some Gorenstein criteria involving the condition that the dual of certain modules are of finite Gorenstein dimension.
1.9 Theorem (See Propositions 6.1, 6.7 and 6.9, and Theorem 6.5).
Suppose that $(R,\mathfrak{m})$ is local. Then $R$ is Gorenstein in each one of the following cases:
@@
    $R$ is a Cohen-Macaulay local ring, and there exists an $R$-module $M$ such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and that satisfies any one of the following:
@@
    $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ for all finitely generated $R$-modules $M$.
@@
    $R$ has positive depth, and every finitely generated $R$-module $M$ whose dual $M^{\ast}$ is non-zero and of finite Gorenstein dimension, is also of finite Gorenstein dimension.
@@
In the last part of this paper, as an application, we focus on the Kähler differential modules $\Omega_{X/Y}^{(n)}$ and the derivation modules $\mathcal{D}er_{Y}^{n}(X)$, defined over locally Noetherian schemes, where $n\geq 1$ is an integer. Initially, we investigate the conditions under which the Kähler differential module $\Omega_{X/Y}^{(n)}$ is reflexive, $k$-torsionfree, and $k$-syzygy for $k=1,2$ (see Theorem 8.1, and Corollaries 8.2 and 8.3). The analysis of these properties in Kähler differential modules of order $n$, in the context of regularity questions of affine rings, has a long history in the literature, being discussed by several authors, especially in the context of algebraic varieties and analytic spaces, particularly when $Y$ is the spectrum of a field (see, for example, [Kunz, Milher, graf, Lipman, Suzuki, Vetter, Lebelt, Greuel]).
Furthermore, we consider the Herzog-Vasconcelos conjecture, which states: Let $R$ be a local $k$-algebra, where $k$ is a field of characteristic zero. If ${\rm pd}_{R}({\rm Der}_{k}(R))<\infty$, then ${\rm Der}_{k}(R)$ is a free $R$-module. This conjecture remains open, although some particular cases have already been resolved (see [He-M, Section 4], [HV, PerezNeto]). Motivated by this conjecture, we propose a generalized version in the context of schemes, called the generalization of the Herzog-Vasconcelos conjecture (GHVC), which is described as follows: Let $n\geq 1$ and ${\rm pd}_{\mathcal{O}_{X}}{\mathcal{D}er}_{Y}^{n}(X)<\infty$. Under what conditions on $X$ and $Y$, and for which values of $n$, does this imply that ${\mathcal{D}er}_{Y}^{n}(X)$ is locally free? (see Question 8.6). Following this line of investigation, we present an application, in which we provide a partial answer to Question 8.6 (see Proposition 8.7).
@@
The organization of this paper is as follows. In Section 2 we provide definitions, notations, and some results that are considered in this paper. In Section 3 we prove Theorem 1.1 and explore some of its consequences. In Section 4, we prove Theorem 1.3, and in Section 5 we provide freeness criteria including Theorems 1.7 and 1.8. In Section 6, we discuss some questions about the Gorensteiness of a ring involving Gorenstein dimension of the dual of modules, and prove Theorem 1.9. In Section 7, we introduce the notation and review some facts concerning Kähler differential modules and differential modules over rings and schemes. Subsequently, in Section 8, we explore the properties of the Kähler differential module $\Omega_{X/Y}^{(n)}$, focusing on its reflexivity, $k$-torsionfreeness, and $k$-syzygy properties, particularly for $k=1,2$. Finally, in the same section, we offer a partial answer to the generalized Herzog-Vasconcelos conjecture (GHVC).
@@ 2. Setup and preliminaries
Throughout this paper, $R$ is a commutative Noetherian ring, and all $R$-modules are considered to be finitely generated.
@@
Let $M$ and $N$ be $R$-modules, and let $k$ be a non-negative integer.
@@
    Let P:⋯⟶P_i ⟶φi P_i-1 ⟶⋯⟶P_1 ⟶φ1 P_0 ⟶φ0 M ⟶0 be a projective resolution of $M$.
@@
      The Auslander transpose of $M$, denoted by $\operatorname{Tr}(M)$, is defined as the cokernel of the induced map $\varphi_{1}^{\ast}:P_{0}^{\ast}\to P_{1}^{\ast}.$ When $k\geq 1$, we set T_k(M)=Tr(Ω^k-1(M)).
@@
    For $k\geq 1$, $M$ is said to be a $k$-syzygy if there exists an exact sequence 0 →M →P_k-1 →P_k-2 →⋯→P_1 →P_0, where each $P_{i}$ is projective. For $k=0$, every $R$-module is said to be a 0-syzygy.
@@
    We say that $M$ satisfies $(S_{k})$ (resp. $(\widetilde{S}_{k})$) if $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\min\{k,\operatorname{ht}(\mathfrak{p})\}$ (resp. $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\min\{k,0pt_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\})$ for all prime ideals $\mathfrak{p}$ of $R$.
@@
The $R$-modules $\operatorname{Tr}(M)$ and $\Omega^{k}(M)$ ($k\geq 0$) are uniquely determined up to projective summands. It is easy to see that $\operatorname{Tr}(\operatorname{Tr}(M))\approx M$. Note that $M$ is a $k$-syzygy if and only if there exists an $R$-module $L$ such that $M\approx\Omega^{k}(L)$. Whenever $R$ is local, we consider the Auslander transpose and the syzygies (of an $R$-module) defined using minimal free resolutions. In this case, these $R$-modules are defined uniquely up to isomorphism (rather than projective equivalence).
The notion of Gorenstein dimension was introduced by Auslander [AnneauxdeGorensteinettorsionenalgebrecommutative] and developed by Auslander and Bridger in [Stablemoduletheory].
@@
    The Gorenstein dimension of $M$, denoted by $\operatorname{G-dim}_{R}(M)$, is defined to be the infimum of all non-negative integers $k$ such that there exists an exact sequence 0 →G_k →⋯→G_0 →M →0, where each $G_{i}$ is totally reflexive.
@@
    ([GorensteinDimensions, Theorem 1.2.7]) If $M\not=0$ and $\operatorname{G-dim}_{R}(M)<\infty$, then G-dim_R(M)=sup{i≥0: Ext_R^i(M,R)/=0}.
@@
    ([Stablemoduletheory, Lemma 3.19(1)]) $M$ is totally reflexive if and only if $\operatorname{Tr}(M)$ is totally reflexive.
@@
    ([GorensteinDimensions, Corollary 1.2.9]) Let $0\to M\to N\to L\to 0$ be an exact sequence of $R$-modules. If two $R$-modules of the sequence have finite Gorenstein dimension, then so has the third.
@@
    ([GorensteinDimensions, Proposition 1.2.10]) $\operatorname{G-dim}_{R}(M)\leq\operatorname{pd}_{R}(M)$ with equality if $\operatorname{pd}_{R}(M)<\infty$.
@@
    ([GorensteinDimensions, Theorem 1.4.8]) If $M\not=0$ and $\operatorname{G-dim}_{R}(M)<\infty$, then G-dim_R(M)+ depth_R(M)=depth_R(R). In particular, $\operatorname{G-dim}_{R}(M)\leq\operatorname{depth}_{R}(R)$.
@@
    ([GorensteinDimensions, Theorem 1.4.9]) Let $k$ be the residual field of $R$. The following conditions are equivalent:
@@
      $\operatorname{G-dim}_{R}(M)<\infty$ for all (finitely generated) $R$-modules $M$.
@@
The formula given in Facts 2.3(5) is known as the Auslander-Bridger formula. From the definition of the Auslander transpose and Facts 2.3(3) we can derive the following remark.
@@
Let $R$ be a ring and let $M$ be an $R$-module. Then $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ if and only if $\operatorname{G-dim}_{R}(\operatorname{Tr}(M))<\infty$.
3. A generalization of a theorem of Auslander and Bridger
The main goal of this section is to provide a generalization of a celebrated result that Evans and Griffith attribute Auslander and Bridger in [Syzygies, Theorem 3.8] (see also [AscentOfFiniteCohen–MacaulayType, Lemma 1.3]), which characterizes a module to be $k$-torsionfree, $k$-syzygy, involving the condition $(\widetilde{S}_{k})$.